OpenStudy (anonymous):

Convert the function into standard form. y-2=-3(x+4)^2.

OpenStudy (jdoe0001):

\(\bf y-2=-3(x+4)^2\implies y-2=-3(x+4)(x+4)\) expand and rearrage http://www.mathsisfun.com/algebra/standard-form.html